A sample of gas occupies 2.06 L at 25.0 °C and 646 torr. What is the temperature (in °C) in a 275.5 mL container where the pressure is 1.24 atm?
Answer #1

P1V1/T1 = P2V2/T2

P1 = 646 torr = 0.85 atm

V1 = 2.06 L

T1 = 273 + 25 = 298 K

P2 = 1.24 atm

V2 =275.5 mL = 0.2755 L

T2 = ?

(0.85*2.06) / 298 = (1.24*0.2755)/T2

1.75 / 298 = 0.342 / T2

0.00587 = 0.342 / T2

T2 = 0.342 / 0.00587 = 58.2 K

T2 = - 214.8 0C